OpenStudy (anonymous):

REALLY NEED HELP. Consider the line that passes through the point (3, -2) and has a slope of 2. Part 1: Write the equation of this line using point-slope form. Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

So point-slope is

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

I know the formulas..

OpenStudy (anonymous):

y+2=2(x-3)

OpenStudy (anonymous):

For part 1

OpenStudy (anonymous):

To write this in slope intercept

OpenStudy (anonymous):

Slope intercept is y=mx+b

OpenStudy (anonymous):

Which is going to be

OpenStudy (anonymous):

Also for part 1 it is y+2=3(x-3)

OpenStudy (anonymous):

For Part 2

OpenStudy (anonymous):

Gah

OpenStudy (anonymous):

It was 2

OpenStudy (anonymous):

y+2=2(x-3)

OpenStudy (anonymous):

y=2x-2

OpenStudy (anonymous):

For part 2

OpenStudy (anonymous):

and standard form is ax+by=c

OpenStudy (anonymous):

To get that 3x+(-2)=2

OpenStudy (anonymous):

Verify yourself to check

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

How to solve

OpenStudy (anonymous):

ok, so how would I prove the work for example, from point slope intercept form to slope intercept?